OpenStudy (anonymous):

this problems really confusing me, (1)/(5^x)=5^6+x how would you solve for x?

OpenStudy (kinggeorge):

I'm assuming you mean\[\frac{1}{5^x}=5^{6+x}\]correct?

OpenStudy (kinggeorge):

If that's the case, then notice that \[\frac{1}{5^x}=5^{-x}\]by definition.

OpenStudy (kinggeorge):

Hence, you can equate the exponents to get \(-x=6+x\). Can you finish the problem from here?

OpenStudy (anonymous):

well that's not how the problem was typed, but maybe they made a typo. but yes thank you I can finish it from there

OpenStudy (kinggeorge):

Was it instead typed as \[\frac{1}{5^x}=5^{6}+x\]